# Real-to-Sim Robot Policy Evaluation with Gaussian Splatting Simulation of Soft-Body Interactions

### v1

@@ I Introduction
Robotic manipulation policies have advanced rapidly across a wide range of tasks [chi2023diffusionpolicy, zhao2023learning, chi2024universal, lin2025sim, tang2023industreal]. However, their evaluation still relies heavily on real-world trials, which are slow, expensive, and difficult to reproduce. As the community shifts toward training foundation models for robotics [rt22023arxiv, black2024pi0, intelligence2025pi05, gr00tn1_2025, trilbmteam2025careful, geminiroboticsteam2025geminirobotics], whose development depends on rapid iteration and large-scale benchmarking, this reliance has become a significant bottleneck.
@@ II-A Robot Policy Evaluation
Evaluating robot policies is essential for understanding and comparing policy behaviors. Most systems are still evaluated directly in the real world [trilbmteam2025careful, octo_2023, pi0experimentwild, padalkar2023open, khazatsky2024droid], but such evaluations are costly, time-consuming, and usually tailored to specific tasks, embodiments, and sensor setups. To enable more systematic study, prior works have introduced benchmarks, either in the real world through standardized hardware setups [Calli_2015, van2016robotic, 7583659, zhou2023trainofflinetestonline, dasari2022rb2roboticmanipulationbenchmarking] or in simulation through curated assets and task suites [taomaniskill3, james2019rlbenchrobotlearningbenchmark, li2024behavior1k, pmlr-v164-srivastava22a, puig2023habitat30cohabitathumans, robocasa2024, zhu2025robosuitemodularsimulationframework, mandlekar2023mimicgendatagenerationscalable, yang2025robotpolicyevaluationsimtoreal, 7583659, wang2025roboevalroboticmanipulationmeets]. Real-world benchmarks offer high fidelity but lack flexibility and scalability, while simulators often suffer from unrealistic dynamics and rendering, which limits their reliability as proxies for physical experiments. This is widely referred to as the “sim-to-real gap” [peng2018sim, chebotar2019closing, openai2019solvingrubikscuberobot, ho2021retinaganobjectawareapproachsimtoreal]. We aim to narrow this gap by building a realistic simulator that combines high-quality rendering with faithful soft-body dynamics. Compared to SIMPLER [li2024evaluating], which relies on green-screen compositing, and Real-is-sim [abouchakra2025realissim], which focuses on rigid-body simulation, our method integrates Gaussian Splatting-based rendering with soft-body digital twins derived from interaction videos, eliminating the dependence on static cameras and providing more realistic appearance and dynamics.
@@ IV-A2 Evaluation
We evaluate four state-of-the-art imitation learning policies: ACT [zhao2023learning], DP [chi2023diffusionpolicy], SmolVLA [shukor2025smolvla], and Pi-0 [black2024pi0]. The real-world setup consists of a single UFactory xArm 7 robot arm equipped with two calibrated Intel RealSense RGB-D cameras: a D405 mounted on the robot wrist and a D455 mounted on the table as a fixed external camera. All policies take as input images from both camera views, along with the current end-effector state. For push-T, the end-effector state includes only the 2D position $(x,y)$; for the other tasks, it additionally includes the position, rotation, and gripper openness. Across all tasks, we collect 39-60 successful demonstrations via teleoperation using GELLO [wu2023gello]. Training is performed using the open-source LeRobot [cadene2024lerobot] implementation, except for Pi-0, where we adopt the original implementation [black2024pi0] for better performance.
@@ IV-D Policy Performance Analysis
IsaacLab [nvidia2024isaac] - - 0.022 0.237 0.031 0.649
Ours w/o color 0.200 0.805 0.156 0.714 0.031 0.529
Ours w/o phys. 0.200 0.694 0.119 0.832 0.031 0.905
Ours 0.087 0.944 0.096 0.901 0.000 0.915
@@ A-D2 Episode Settings
In all evaluation experiments in the main paper, the number of episodes for each task and the grid-based initial configuration randomization ranges are set as in Table VII.

### v2

@@ I Introduction
Robotic manipulation policies have advanced rapidly across a wide range of tasks [chi2023diffusionpolicy, zhao2023learning, chi2024universal, lin2025sim, tang2023industreal]. However, their evaluation still relies heavily on real-world trials, which are slow, expensive, and difficult to reproduce. As the community shifts toward training foundation models for robotics [rt22023arxiv, kim2024openvla, black2024pi0, intelligence2025pi05, gr00tn1_2025, trilbmteam2025careful, geminiroboticsteam2025geminirobotics], whose development depends on rapid iteration and large-scale benchmarking, this reliance has become a significant bottleneck.
@@ II-A Robot Policy Evaluation
Evaluating robot policies is essential for understanding and comparing policy behaviors. Most systems are still evaluated directly in the real world [trilbmteam2025careful, octo_2023, pi0experimentwild, padalkar2023open, khazatsky2024droid, zhou2025autoeval, atreya2025roboarena], but such evaluations are costly, time-consuming, and usually tailored to specific tasks, embodiments, and sensor setups. To enable more systematic study, prior works have introduced benchmarks, either in the real world through standardized hardware setups [Calli_2015, van2016robotic, 7583659, zhou2023trainofflinetestonline, dasari2022rb2roboticmanipulationbenchmarking] or in simulation through curated assets and task suites [taomaniskill3, james2019rlbenchrobotlearningbenchmark, li2024behavior1k, pmlr-v164-srivastava22a, puig2023habitat30cohabitathumans, robocasa2024, zhu2025robosuitemodularsimulationframework, mandlekar2023mimicgendatagenerationscalable, yang2025robotpolicyevaluationsimtoreal, wang2025roboevalroboticmanipulationmeets, badithela2025reliable]. Real-world benchmarks offer high fidelity but lack flexibility and scalability, while simulators often suffer from unrealistic dynamics and rendering, which limits their reliability as proxies for physical experiments. This is widely referred to as the “sim-to-real gap” [peng2018sim, chebotar2019closing, openai2019solvingrubikscuberobot, ho2021retinaganobjectawareapproachsimtoreal]. We aim to narrow this gap by building a realistic simulator that combines high-quality rendering with faithful soft-body dynamics. Compared to SIMPLER [li2024evaluating] and RobotArena $\infty$ [jangir2025robotarena] which relies on green-screen compositing, Ctrl-World [guo2025ctrl] which relies on 2D video world models, and Real-is-sim [abouchakra2025realissim] which focuses on rigid-body simulation, our method integrates Gaussian Splatting-based rendering with soft-body digital twins derived from interaction videos, eliminating the dependence on static cameras and providing more realistic appearance and dynamics.
@@ IV-A2 Evaluation
We evaluate four state-of-the-art imitation learning policies: ACT [zhao2023learning], DP [chi2023diffusionpolicy], Pi-0 [black2024pi0], and SmolVLA [shukor2025smolvla]. The real-world setup consists of a single UFactory xArm 7 robot arm equipped with two calibrated Intel RealSense RGB-D cameras: a D405 mounted on the robot wrist and a D455 mounted on the table as a fixed external camera. All policies take as input images from both camera views, along with the current end-effector state. For push-T, the end-effector state includes only the 2D position $(x,y)$; for the other tasks, it additionally includes the position, rotation, and gripper openness. Across all tasks, we collect 39-60 successful demonstrations via teleoperation using GELLO [wu2023gello]. Training is performed using the open-source LeRobot [cadene2024lerobot] implementation, except for Pi-0, where we adopt the original implementation [black2024pi0] for better performance.
@@ IV-D Policy Performance Analysis
IsaacLab [nvidia2024isaac] - - 0.270 0.237 0.196 0.649
Ours w/o color 0.200 0.805 0.343 0.714 0.354 0.529
Ours w/o phys. 0.241 0.694 0.248 0.832 0.100 0.905
Ours 0.076 0.944 0.174 0.901 0.108 0.915
@@ A-D2 Episode Settings
In all evaluation experiments in the main paper, the number of episodes for each task and the grid-based object initial configuration randomization ranges (relative to their prespecified poses) are set as in Table VII.